OpenStudy (anonymous):

a boat leaves the entrance of a harbor and travels 83 miles on a bearing on N19E. How many miles north and how many miles east from the the harbor has the boat traveled?

OpenStudy (jdoe0001):



OpenStudy (anonymous):

ok i got it is to multiply 83 times cos(19) it give us x = 78.5 and then we get y thanks

OpenStudy (jdoe0001):

well, you want to know how far East the boat went the boat traveled N19E on North Bearing, that means as you see in the picture 19 degrees from the North-South line so you can always find the East distance, the adjacent to the triangle by using \(\bf cos(71^o) = \cfrac{adj}{83} \implies cos(71^o) \times 83 = adj\)

OpenStudy (jdoe0001):

or using the 19 degrees angle instead, it'd be \(\bf sin(19^o) = \cfrac{opp}{83} \implies sin(19^o) \times 83 = opp\)

OpenStudy (anonymous):

I did in a different way and i got the same results but that's a lot more easier.

OpenStudy (jdoe0001):

tis ok :)

OpenStudy (anonymous):

thanks so much the way you did it will save me some time when i do the exam ;)

OpenStudy (jdoe0001):

yw